OpenStudy (anonymous):

1,440 is 120 percent of _________.

OpenStudy (anonymous):

1440 +288 = 1748

OpenStudy (anonymous):

120/100*n=1440 then solve for n

OpenStudy (anonymous):

1440/1.2=1200